# Ant Colony Sampling with GFlowNets for Combinatorial Optimization

### v2

@@ 6.2 Comparison with other reinforcement learning methods for combinatorial optimization
Evaluation on real-world datasets. We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib [64] and CVRPLib [72] X datasets. The models were trained on random uniform instances of TSP/CVRP with sizes 200 for dataset sizes 100-299, 500 for sizes 300-699, and 1000 for sizes 700-1499, with all other settings remaining the same. Table 2 presents the results. GFACS outperforms the baselines in TSPLib and large CVRPLib instances but slightly underperforms compared to vanilla ACO on small CVRPLib instances, where ACO with Swap* local search suffices for small-scale CVRP. As the problem scale increases, ACO’s performance declines, underscoring the importance of effective priors using GFlowNets for efficient pheromone posterior updates in large combinatorial spaces.
@@ 6.3 Ablation study
Table 3: Results of the ablation study for the proposed training techniques.
@@
   – Shared $\varepsilon$-norm 17.52 65.51 90.73
    + VarGrad [82] 16.86 64.28 51.95
@@
This section reports the ablation study results, validating our proposed training techniques for stabilizing and enhancing GFlowNets in large-scale combinatorial optimization tasks. As detailed in Table 3, our off-policy training method, incorporating local search, energy reshaping to value potential solutions, and shared $\epsilon$-norm to stabilize conditional GFlowNets training, significantly improves performance across three tasks: TSP, CVRP, and BPP. For comprehensive results, please refer to Table 10.
6.4 Comparative analysis with previous GFlowNets training methods
@@
While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature (e.g., TSP, CVRP) [1], we also compare existing GFlowNets training methods within this domain.
Figure 4: Validation costs during training with different GFlowNet losses.
@@
Forward-looking detailed balance (FL-DB) [59, 84]. FL-DB leverages local credits when the application benefits from useful local rewards for partial solutions. While graph combinatorial optimization tasks are generally locally decomposable [1], yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment. As shown in Figure 4, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G};\theta)$. While FL-DB did not surpass on-policy TB for TSP, our off-policy method demonstrated significant improvement. This trend holds across different deep learning models and inference tasks beyond ACO (see Section D.1).
@@
Gradient of log-parition variance (VarGrad) [65, 82]. The VarGrad loss, initially proposed by Richter et al. [65] for variational inference, was applied to conditional GFlowNets for solving scheduling problems by Zhang et al. [82]. This approach is crucial for combinatorial optimization tasks that require conditional inference on problem instances ($\mathcal{G}$). The key idea involves using a generative distribution and energy to compute the log-partition function through the consistency property of log-partitions over trajectories conditioned on identical instances ($\mathcal{G}$), thereby bypassing the explicit training of the (conditional) log-partition ($\log Z(\mathcal{G})$). Explicit training of the log-partition is challenging in conditional GFlowNets due to the varying energy landscapes across different instance conditions ($\mathcal{G}$). On the other hand, our method, which utilizes explicit training of the log-partition (see Equation 9), outperforms VarGrad in training conditional GFlowNets (see Table 3). This improvement is achieved because our Share $\epsilon$-norms enable stable and explicit training of the log-partition by normalizing energy values into sample means, resulting in an energy distribution across different instances that maintains a mean of zero. Note that accurately estimated the explicit log-partition value is known to stabilize the amortized inference process in GFlowNets [53].
7 Discussion
This paper proposes a novel probabilistic search method leveraging the off-policy training capabilities of GFlowNets. This strategy can be applied not only to ant-colony optimization (ACO) but also to other probabilistic methods requiring an effective learned prior (see Section D.4). Our off-policy framework enables the integration of learning algorithms with offline datasets containing optimal combinatorial solutions for specific problem instances, combined with online exploration to fine-tune the policy through off-to-on training. This approach holds promise for future work.
Limitation. A limitation of our method is that its performance depends on the effectiveness of the local search. A promising future direction could involve leveraging advanced neural-guided improvement heuristics, including neural local search approaches [48, 50], to create a hybrid algorithm that combines GFlowNets with neural improvement-based local search.
@@ A.1 Learning architecture and optimizer
We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by [29] and adopted by [79], as the foundational GNN architecture for the construction of a heuristic matrix. This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks [73] or Graph Convolutional Networks [40].
@@
where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN[25] denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$. In this paper we use $\alpha$ is SiLU[17] and $\mathcal{A}$ is mean pooling. We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$. The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs. The training process is optimized using the AdamW optimizer[46] in conjunction with a Cosine Annealing Scheduler[45]
@@ A.2 Hyperparameters
For TSP and CVRP, where we have local search operators, we also need $\alpha$ for energy reshaping (Eq. 7). We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all TSP and CVRP experiments.
@@ Appendix B Problem settings
In this section, we provide an overview of the combinatorial optimization problems we address. These encompass vehicle routing problems, bin packing problems, and scheduling problems. For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al. [79]. Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^3] https://github.com/henry-yeh/DeepACO from Ye et al. [79].
@@
A CVRPTW instance consists of a set of customer nodes, a single depot node, the demands of each customer node, the vehicle capacity (as in CVRP), and additionally, the earliest and latest arrival times for each node, as well as the service time at each node. We generate the CVRPTW instance following the method provided by Berto et al. [6], which can be accessed here [^4] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py.
@@
An OP instance comprises a set of nodes to visit, each with its position and prize, a depot, and a maximum length constraint. We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$. And we set the prize for each node, following [19, 79], to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$. We set the maximum length to 4 in the OP100 task, following Kool et al. [41].
@@
The bin packing problem (BPP) aims to minimize the number of bins when objects with different volumes need to be packed into a finite number of bins. The cost functions are often modified to maximize the utility of used bins [18, 79]. However, the direct use of the objective function employed in DeepACO can lead to an excessively large $\beta$ since the objective values tend to be saturated near 1.0; note that the objective values can not exceed 1.0 in that setting. Therefore, we slightly modify them to minimize a weighted sum of the inverse of the utilization of the used bins throughout all our BPP experiments.
@@ Appendix C Implementation of destroy-and-repair local search
As demonstrated in Section 5, our algorithm utilizes the local search for off-policy training. The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Eq. 7). For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* [74] and granular neighborhood search from PyVRP [77].
While domain-specific local search techniques are well developed for many combinatorial optimization problems, there are still tasks that lack effective local search solvers like Swap*. A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-reconstruct local search operator. This method is inspired by the large neighborhood search approach described by Pisinger and Ropke [61]. Our implementation utilizes an evolutionary-style algorithm that repeatedly performs destruction, reconstruction, and top-K selection over a specified number of rounds.
@@
Reconstruction with temperature annealing. Given a partial solution, we reconstruct a new solution $x^{\prime}$ using the policy similar to the decoding steps explained in Section 4.2. Considering the purpose of the local search (refining a given solution), we further promote the exploitation by introducing an inverse temperature for the policy. Formally, we modify Equation 5 with the inverse temperature $\delta_{r}$ as follows:
@@
Top-K selection. For a given solution $x$, we perform the aforementioned destruction and reconstruction multiple times (in batch), obtaining a set of refined solutions. Then we evaluate them and leave only top-K solutions among them in terms of solution quality. The remaining solutions are used as an initial solution for the next round of the search.
@@ D.1 Comparison between trajectory balance and forward-looking detailed balance
We trained Attention Model (AM) [41] and heatmap-based policy from Joshi et al. [29] in an unsupervised way using FL-DB and TB loss. Table 6 shows the test results for 128 random TSP50 instances. The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
Table 6: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al. [29] as backbone policies.
@@ D.3 Evaluation with other ACO variants
Table 7: Evaluation results with different pheromone update rules, Elitist AS and MAX-MIN AS. We run $T=10$ ACO rounds with 100 ants in all settings.
@@
AS 10.91 1.77 28.45 1.86
DeepACO + AS 10.77 0.45 28.44 1.80
GFACS + AS 10.75 0.26 28.32 1.40
Elitist AS 10.97 2.31 28.46 1.90
DeepACO + Elitist AS 10.77 0.49 28.45 1.84
GFACS + Elitist AS 10.75 0.28 28.34 1.47
MAX-MIN AS 10.97 2.34 28.46 1.88
DeepACO + MAX-MIN AS 10.78 0.57 28.44 1.82
GFACS + MAX-MIN AS 10.75 0.32 28.33 1.43
Our method is focused on enhancing the design of the heuristic matrix within the ACO framework. Thus, the advancement of the ACO algorithm itself (usually in pheromone update and action selection rule) is orthogonal to our work. To validate this claim, we evaluate our method with different ACO variants other than AS, namely Elitist Ant System [14], MAX-MIN Ant System [68]. The results in Table 7 demonstrate that GFACS consistently outperforms both ACO with human-designed heuristic and DeepACO under all pheromone update rules tested. This underscores the superiority of GFACS in generating a high-quality heuristic matrix.
D.4 Applying our method to other search algorithms
Table 8: Evaluation results combine with active search as posterior inference method.
TSP200 CVRP200
Method Obj. Gap (%) Obj. Gap (%)
Concorde / PyVRP 10.72 0.00 27.93 0.00
Pretrained GFACS
   – Off-policy TB Active Search 10.75 0.29 28.34 1.47
   – On-policy RL Active Search 10.76 0.40 28.38 1.62
Pretrained DeepACO
   – On-policy RL Active Search 10.79 0.65 28.43 1.79
In this subsection, we study active search [3] rather than the pheromone update of ACO. In ACO, pheromone update is a posterior inference given a prior distribution (heuristic) for the probability model that generates the solutions. Similarly, we can think of active search as a posterior update that implicitly updates a given probability model by adjusting the parameters that generate it. We sample 20 solutions and take 5 gradient steps to update the parameter at each iteration. After 50 iterations, we report the average performance of the active search using our method with 3 seeds of 128 instances for each task. Table 8 shows that GAFCS integrates better with active search than DeepACO. This indicates that GAFCS is effective not only with heuristic-based posterior inference but also with learning-based posterior inference.
D.5 Comparison against TAM
Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al. [24], and the result for DeepACO is drawn from Ye et al. [79].

### v4

@@ 5.1 Comprison with different priors
Evaluation on real-world datasets. We compare the performance of our model against baseline models on real-world TSP and CVRP instances from the TSPLib (Reinelt,, 1991) and CVRPLib X-instances (Uchoa et al.,, 2017). The models were trained on random uniform instances of TSP/CVRP with sizes 200 for dataset sizes 100-299, 500 for sizes 300-699, and 1000 for sizes 700-1499, with all other settings remaining the same. Table 1 presents the results. GFACS outperforms the baselines in TSPLib and large CVRPLib instances but slightly underperforms compared to vanilla ACO on small CVRPLib instances, where ACO with Swap* local search suffices for small-scale CVRP. As the problem scale increases, ACO’s performance declines, underscoring the importance of effective priors using GFlowNets for efficient pheromone posterior updates in large combinatorial spaces.
5.2 Comparison with RL solvers
In this section, we present experimental results comparing our method with problem-specific RL approaches that leveraged specialized neural architectures or learning heuristics for the TSP and CVRP.
Baselines. For TSP baselines, we select the AM (Kool et al.,, 2019), POMO (Kwon et al.,, 2020), Sym-NCO (Kim et al.,, 2022) and Pointerformer (Jin et al.,, 2023) as RL-based constructive methods, DIMES (Qiu et al.,, 2022) as RL-based heatmap-based methods, the Select and Optimize (SO; Cheng et al.,, 2023) as RL-based improvement methods. We also applied two search algorithms, the Efficient Active Search (EAS; Hottung et al.,, 2022) and Simulation Guided Beam Search (SGBS; Choo et al.,, 2022), to POMO for a more thorough comparison. For CVRP, [^1] Due to the inaccessibility of source codes, we compare ours with TAM using the reported result in the paper (Hou et al.,, 2023). See Section D.4. we employ POMO and Sym-NCO as baselines with EAS and SGBS. Note that we restrict the duration of these searches to no more than 5 and 10 minutes for problems with 500 and 1000 nodes, respectively. We use Concorde (Applegate et al.,, 2006) for TSP and PyVRP (Wouda et al.,, 2024) for CVRP as oracle heuristic solvers to obtain near-optimal solutions that serve as baselines with $0.00\%$ gap. For ACO algorithms, we use $K=100$ ants with $T=10$ with local search algorithms similar to Ye et al., (2023).
Results. GFACS consistently shows superior or highly competitive performance compared to evaluated baseline methods (Table 2). Ours outperforms fast zero-shot methods like AM and POMO—even when they employ search techniques such as EAS or SGBS—and achieves better results than similarly paced algorithms like DIMES and Pointformer. Although it performs slightly worse than the heuristic solver LKH-3 (Helsgaun,, 2017), GFACS offers faster inference times, especially in CVRP. As a general-purpose algorithm, it holds significant appeal in the competitive benchmarks of both TSP and CVRP. Notably, despite its small neural architecture and requiring less than three hours of training, GFACS outperforms larger models like POMO and Sym-NCO, which need over a week of training (Kim et al.,, 2022).
@@ 5.3 Ablation study
Table 3: Results of the ablation study.
@@
   – Shared Energy Norm. 17.52 65.51 90.73
    + VarGrad 16.86 64.28 51.95
@@
This section reports the ablation study results, validating our proposed training techniques for stabilizing and enhancing GFlowNets in large-scale combinatorial optimization tasks. As detailed in Table 3, our off-policy training method, incorporating local search, energy reshaping to value potential solutions, and shared $\epsilon$-norm to stabilize conditional GFlowNets training, significantly improves performance across three tasks: TSP, CVRP, and BPP. For comprehensive results, including standard deviations, refer to Section D.5.
5.4 Comparison with other GFlowNet-based training methods
@@
While this paper proposes a novel training method for GFlowNets designed for large-scale combinatorial optimization problems that are not locally decomposable due to their cyclic nature, such as in TSP and CVRP, we also compare existing GFlowNets training methods within this domain.
Figure 5: Validation cost on TSP with 200 nodes during training, compared to forward-looking detailed balance (FL-DB).
@@
Forward-looking detailed balance (FL-DB; Pan et al.,, 2023; Zhang et al.,, 2023). FL-DB leverages local credits when the application benefits from useful local rewards for partial solutions. While graph combinatorial optimization tasks are generally locally decomposable (Ahn et al.,, 2020), yielding useful local rewards, tasks like TSP do not offer substantial local rewards due to their cyclic nature, which imposes a global constraint, thus complicating local credit assignment. As shown in Figure 5, various GFlowNet methods were used to train the ACO’s heuristic matrix $\eta(\mathcal{G})$. While FL-DB did not surpass on-policy TB for TSP, our off-policy method demonstrated significant improvement. This trend holds across different deep learning models and inference tasks beyond ACO; see Section D.1.
@@
Gradient of log-partition variance (VarGrad). Originally proposed by Richter et al., (2020) for variational inference, the VarGrad loss was applied to conditional GFlowNets for scheduling problems by Zhang et al., 2022b . This approach is vital for combinatorial optimization tasks requiring conditional inference on the problem instance $\mathcal{G}$. It computes the log-partition function via the consistency property over trajectories conditioned on identical instances, thus bypassing explicit training of $\log Z(\mathcal{G})$. Explicit training is challenging due to varying energy landscapes across different instances. Our method, utilizing explicit training of the log-partition $\log Z(\mathcal{G})$ as in Eq.(8), outperforms VarGrad in training conditional GFlowNets (Table 3). This improvement stems from our Shared energy normalization, which stabilizes training by normalizing energy values to sample means, resulting in an energy distribution with a zero mean across instances. Note that accurately estimating the explicit log-partition value is known to stabilize training of GFlowNets (Malkin et al.,, 2023).
5.5 Integrating GFACS with active search
Background. In this section, we explore the integration of active search (AS) as a posterior update on a parameterized prior distribution $p_{\theta}(\cdot|\mathcal{G})$. In AS, the prior parameter $\theta$, pretrained on a large set of random problem instances $\mathcal{G}$, is adaptively updated for specific test-time instances $\mathcal{G}_{\text{test}}$ using the same objective function as in the prior training. We trained this prior $p_{\theta}$ using two methods: on-policy RL method (Kwon et al.,, 2020) that employ REINFORCE with shared baselines, and our off-policy GFlowNet training method as explained in Section 4.
Setting. For the active search procedure, we sample 20 solutions per test-time instance and perform 5 gradient updates at each iteration, updating $\theta$ over a total of 50 iterations.
Table 4: Evaluation results combine with active search (AS) as a posterior search method.
TSP200 CVRP200
Method Obj. Gap (%) Obj. Gap (%)
Concorde / PyVRP 10.72 0.00 27.93 0.00
RL prior + AS 10.79 $\pm$ 0.00 0.65 28.43 $\pm$ 0.02 1.79
GFlowNet prior + AS 10.75 $\pm$ 0.00 0.29 28.34 $\pm$ 0.01 1.47
Results. As shown in Table 4, the off-policy GFlowNet prior combined with AS posterior inference outperforms the RL counterpart. This result provides further evidence supporting our hypothesis that the multi-modal prior from GFlowNet training synergizes effectively with the subsequent iterative posterior update algorithm. Additionally, the results demonstrate the flexibility of our method beyond the ACO, suggesting a promising direction for future research into the benefits of using a multi-modal informative prior.
6 Conclusion
In this paper, we introduced a novel probabilistic meta-heuristic that hierarchically combines GFlowNets-based amortized multi-modal priors with posterior search using ant colony optimization (ACO) for combinatorial optimization problems. Our strategy achieved remarkable success across seven benchmark combinatorial optimization tasks. Furthermore, by applying our meta-heuristic to other posterior search methods like active search (AS), we demonstrated the expandability and modularity of our approach. Limitation of our work is that, while our method empirically improves solution quality, we do not provide formal theoretical guarantees in how our method gives better optimality than the reinforcement learning counterpart.
Acknowledgment
The authors thank Haoran Ye and Federico Berto for valuable discussions for this project. This research is based on funding from Samsung, Intel, CIFAR and the CIFAR AI Chair program. The research was enabled in part by computational resources provided by the Digital Research Alliance of Canada (https://alliancecan.ca), Mila (https://mila.quebec), and NVIDIA. This work was also partially supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2024-00410082).
@@ A.1 Learning architecture and optimizer
We utilize the anisotropic graph neural network with edge gating mechanisms, which is proposed by Joshi et al., (2021) and adopted by Ye et al., (2023), as the foundational GNN architecture for the construction of a heuristic matrix. This network offers node and edge embeddings that are more beneficial for edge prediction than other popular GNN variants, such as Graph Attention Networks (Veličković et al.,, 2018) or Graph Convolutional Networks (Kipf and Welling,, 2017).
@@
where $U^{\ell},V^{\ell},P^{\ell},Q^{\ell},R^{\ell}\in R^{d\times d}$ are the learnable parameters of layer $\ell$, ACT denotes the activation function, BN (Ioffe and Szegedy,, 2015) denotes batch normalization, $\mathcal{A}$ denote the aggregation function, $\sigma$ is the sigmoid function, $\odot$ is Hadamard product, $\mathcal{N}_{i}$ denotes the neighborhoods of node $v_{i}$. In this paper we use ACT is SiLU (Elfwing et al.,, 2018) and $\mathcal{A}$ is mean pooling. We transform the extracted edge features into real-valued heuristic metrics by employing a 3-layer Multilayer Perceptron (MLP) that includes skipping connections to the embedded node feature $h$. The SiLU is the activation function for all layers except the output layer, in which the sigmoid function is applied to generate normalized outputs. The training process is optimized using the AdamW optimizer (Loshchilov and Hutter, 2017a, ) in conjunction with a Cosine Annealing Scheduler (Loshchilov and Hutter, 2017b, )
@@ A.2 Hyperparameters
In Equation 7, we also need a hyperparameter $\alpha$ that controls how much we augment the energy of a solution with the energy of the refined solution. If $\alpha=0$, we do not augment at all, and if $\alpha=1$, the energy of a solution is defined to be equal to the energy of the refined solution. We empirically found that linearly increasing $\alpha$ from $0.5$ to $1.0$ along the epochs works well, and we use that setting for all experiments.
@@ Appendix B Problem settings
In this section, we provide an overview of the combinatorial optimization problems we address. These encompass vehicle routing problems, bin packing problems, and scheduling problems. For a comprehensive understanding of each problem’s specific settings, we direct readers to the work by Ye et al., (2023). Furthermore, for access to the exact source code pertaining to each problem, please refer to the repository available at  [^2] https://github.com/henry-yeh/DeepACO from Ye et al., (2023).
@@
A CVRPTW instance consists of a set of customer nodes, a single depot node, the demands of each customer node, the vehicle capacity (as in CVRP), and additionally, the earliest and latest arrival times for each node, as well as the service time at each node. We generate the CVRPTW instance following the method provided by Berto et al., (2023), which can be accessed here. [^3] https://github.com/ai4co/rl4co/blob/main/rl4co/envs/routing/cvrptw/generator.py
@@
An OP instance comprises a set of nodes to visit, each with its position and prize, a depot, and a maximum length constraint. We sample a random OP instance by first sampling the position of nodes, including the starting node, from a unit square $[0,1]^{2}$. And we set the prize for each node, following Fischetti et al., (1998); Ye et al., (2023), to be $p_{i}=(1+\lfloor 99\cdot\frac{d_{0,i}}{\max_{j=1}^{n}{d_{0,j}}}\rfloor)/100$, where $d_{0,i}$ is the Euclidean distance from the depot to node $i$. We set the maximum length to 4 in the OP100 task, following Kool et al., (2019).
@@
The bin packing problem (BPP) aims to minimize the number of bins when objects with different volumes need to be packed into a finite number of bins. The cost functions are often modified to maximize the utility of used bins (Falkenauer and Delchambre,, 1992; Ye et al.,, 2023). However, the direct use of the objective function employed in DeepACO can lead to an excessively large $\beta$ since the objective values tend to be saturated near 1.0; note that the objective values can not exceed 1.0 in that setting. Therefore, we slightly modify them to minimize a weighted sum of the inverse of the utilization of the used bins throughout all our BPP experiments.
@@ Appendix C Implementation of destroy-and-repair local search
As demonstrated in Section 4, our algorithm utilizes the local search for off-policy training. The refined solutions from local search serve as exploitation batches, and they are also used for energy-reshaping (Equation 7). For some problems like traveling salesman problem (TSP), capacitated vehicle routing problem (CVRP), and CVRP with time window (CVRPTW), we used the well-established local search operators such as 2-opt, Swap* (Vidal,, 2022) and granular neighborhood search from PyVRP (Wouda et al.,, 2024).
While domain-specific local search techniques are well developed for many combinatorial optimization problems, there are still tasks that lack effective local search solvers like Swap*. A potential candidate for a general-purpose local search solver in combinatorial optimization is the destroy-and-repair local search operator. This method is inspired by the large neighborhood search approach described by Pisinger and Ropke, (2019). Our implementation utilizes an evolutionary-style algorithm that repeatedly performs destruction, reconstruction, and top-K selection over a specified number of rounds.
@@
Repair with temperature annealing. Given a partial solution, we reconstruct a new solution $x^{\prime}$ using the policy similar to the decoding steps explained in Section 3.2. Considering the purpose of the local search (refining a given solution), we further promote the exploitation by introducing an inverse temperature for the policy. Formally, we modify Equation 5 with the inverse temperature $\delta_{r}$ as follows:
@@
Top-K selection. For a given solution $x$, we perform the aforementioned destruction and repair multiple times (in batch), obtaining a set of refined solutions. Then we evaluate them and leave only top-K solutions among them in terms of solution quality. The remaining solutions are used as an initial solution for the next round of the search.
@@ D.1 Comparison between trajectory balance and forward-looking detailed balance
Table 7: Comparison between forward-looking detailed balance (FL-DB) and trajectory balance (TB) loss in TSP50, using Attention Model (AM) and heatmap-based method from Joshi et al., (2021) as backbone policies.
@@
We trained Attention Model (AM; Kool et al.,, 2019) and heatmap-based policy from Joshi et al., (2021) in an unsupervised way using FL-DB and TB loss. Table 7 shows the test results for 128 random TSP50 instances. The result is aligned with the previous observation from GFACS, showing the limitation of FL-DB loss for solving complex CO tasks.
@@ D.3 Evaluation with other ACO variants
Table 8: Evaluation results with different pheromone update rules, Elitist Ant System (Elitist) and MAX-MIN Ant System (MAX-MIN). We run $T=10$ ACO rounds with 100 ants in all settings.
@@
Ant System 10.91 1.77 28.45 1.86
DeepACO + Ant System 10.77 0.45 28.44 1.80
GFACS + Ant System 10.75 0.26 28.32 1.40
Elitist 10.97 2.31 28.46 1.90
DeepACO + Elitist 10.77 0.49 28.45 1.84
GFACS + Elitist 10.75 0.28 28.34 1.47
MAX-MIN 10.97 2.34 28.46 1.88
DeepACO + MAX-MIN 10.78 0.57 28.44 1.82
GFACS + MAX-MIN 10.75 0.32 28.33 1.43
Our method is focused on enhancing the design of the heuristic matrix within the ACO framework. Thus, the advancement of the ACO algorithm itself (usually in pheromone update and action selection rule) is orthogonal to our work. To validate this claim, we evaluate our method with different ACO variants other than Ant System, namely Elitist Ant System (Dorigo et al.,, 1996), MAX-MIN Ant System (Stützle and Hoos,, 2000). The results in Table 8 demonstrate that GFACS consistently outperforms both ACO with human-designed heuristic and DeepACO under all pheromone update rules tested. This underscores the superiority of GFACS in generating a high-quality heuristic matrix.
D.4 Comparison against TAM
Table 9: Results for CVRP instances used in TAM. The results for AM and TAM is drawn from Hou et al., (2023), and the result for DeepACO is drawn from Ye et al., (2023).